Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day11


Crew #108 Log Book for December 13, 2011

Photos of the day





Commander's Check-In Report

Date and time: 

13 December 2011, 17:20 

Written by : 

Charlotte Poupon, Crew 108, Commander

Crew Physical Status: 

We all are well

Time Departed/Returned from EVA: 

EVA # 19 : from 09:37 to 11:29 

Brief Narrative of Field Mission Results: 

It was a pedestrian excursion. It appears the snow makes walking quite slippery and problematic. 

EVA Data/Interpretations: 

Wet snow on exposed clay sediments exhibit a very low coefficient of kinetic friction (estimated between 0.15 and 0.25). HabCar use is not recommended in these conditions. 

Engineering/Hab Maintenance: Engineers planned couple of repairs and waiting for Mission Support to go on duty. 

Report Transmission Schedule: All submitted.

Plans for Tomorrow: 

Not yet determined Inventory: everything is OK 

Support Requested: 

Engineers ask a GO to fix a breathing apparatus case broken during an EVA (without crew member injury) Engineers ask a GO from Peter to fix the antenna 

 Miscellaneous:

Commander's Report

Date: 

13 December 2011

Written By: 

Charlotte Poupon, Crew 108, Commander

Usha has illuminated our day twice preparing panecakes for breakfast and cookies for afternoon tea.

The EVA morning confirmed difficulties in wandering (including the minor fall from a crew member who broke his breathing bagback) due to snowfalls, so the afternoon was devoted to domestic activities such as thinking about fixing the damaged gear, playing chest, or personal writing.

We received a mail from Robert Zubrin asking for the summary report, due to friday. This exercise makes us aware of our achievments and is pretty  stimulating for the coming days.

Science Report




Engineering Report

Date and time: 

13 December 2011, 17:50

Written by:

Mike Lotto, Crew 108, Engineer

Generator/Electricity:

Did you have any electricity outage? (Yes/No): No

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.375
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): Full
Notes/Comments:

Oil:

ATV Oil Quantity (# of Quarts in storage): 1.25
Car Oil Quantity (# of Quarts in storage): 1.25

Water Status:

Outside Potable Water Tank Level (inches from bottom): 18
Trailer Potable Water Tank Level (inches from bottom): Approximately 30 inches (or 70% full)
Hab Potable Water Tank Level (inches from bottom): 11
Potable Water Meter Reading: Unknown
In to GreenHab Meter Reading: Offline
Notes/Comments:

GreenHab:

Condition of plants in Tank 1: Offline
Condition of plants in Tank 2: Offline
Condition of plants in Tank 3: Offline
Greenhab Notes/Comments: GreenHab is offline this season.

Observatory:

Telescope Used? (Yes/No): No
Observatory Notes/Comments: We've completed the on-line training class, and we request the keys to the observatory.

Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: Good
Odometer at end of day: 125464 miles
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 03:07
Comments:

ATV 2 (Spirit):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 01:11
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 03:07
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 67
Actual temperature upstairs (Farenheit): 69

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 50

Computers/Networking Infrastructure:
Notes/comments on internet/computers?: None

General Engineering Notes/Comments: EVA Suit 4 is disabled at the moment.

Journalist Report

Date and time:

13 December 2011, 19:05

Written by:

Mike Lotto, Crew 108, Engineer

The day started with a relaxed morning as the crew awoke to a snowy landscape. Other than sending several emails to Mission Support, the team waited in the Hab for a warmer outside air temperature. At 9:30 AM and a balmy air temperature of 32 degrees, the day's EVA slowly began on foot due to poor road conditions. After reaching the base of Phobos Peak, the EVA team turned North and eventually climbed Sagewood Peak. While returning to the Hab to complete the EVA, a crew member slipped on a patch of ice, which was treacherously hidden by snow, and damaged their EVA suit's backpack. The remainder of the day consisted of discussions regarding a fix for the backpack and intense games of chess. Ultimately, the crew is truly enjoying their time and can't believe how quickly the days are passing by.



Chef's Report

Date:

13/12/2011 

Written by: Ashley Dale, Crew 108, Engineer

1a) Today was a cooking day.

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

1 x pack of Jambalaya mix (serves 8 – with addition of meats etc)

2 x tinned shrimp

1.5 x cup of dehydrated tomatoes

Egg powder to make the equivalent of 5 eggs

Parsley

Basil

Garlic powder

Olive Oil

3) List any main meal foods not finished by the crew

1/3 of a bowl not finished.

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Scrambled egg:

12 x tbsp of dehydrated egg powder

12 x tbsp of water

2 x tbsp of olive oil

½ x tsp Parsley

½ x tsp Basil

½ x tsp garlic powder

Mix together constantly and heat up slowly. Turn off heat when contents are nearly complete and take off hob. Add parsley, basil, and garlic powder and mix in. Internal heat of content should finish it off.

5) Lessons learned (if any)

6) Comments/questions for Kim and Jean, the food study investigators

Fixed the metal grid over the hobs – good as new!

The jambalaya was a big hit amongst the crew – I’d recommend keeping it available.